# Joint AP-UE Association and Power Factor Optimization for Distributed Massive MIMO

### 22 Feb 2024 v1

@@ III Proposed Joint Solution using Relaxation Technique
The optimization problem (15a) is a mixed-integer type problem which has a very high complexity, so we convert a binary constraint (15b) into a continuous variable, thus the constraint becomes :
$d_{mt}=[0,1],\forall m\in\mathcal{M},\ t\in\mathcal{T},$ (16)
Remark 1 : The introduction of l1-penalty in the objective function serves to induce the sparsity in AP-UE association matrix D. By careful selection of the suitable trade-off factor $\alpha$, a controlled-sparsity is achieved by limiting only few APs to serve the UEs. Therefore, the matrix D has non-zero values only for the chosen AP-UE pair. After completion of AP-UE association, all entries in D undergo rounding to the nearest integer values i.e., zero and one, thus maintaining the feasibility of the optimization problem (15a). In the forthcoming result section, the SE is compared with and without rounding-off of values in D to establish the effectiveness of our proposed approach.
Then, the optimization problem (15a) can be formulate as :
$\underset{\mathbf{\eta^{u}},\textbf{D},\textbf{I}}{\max~{}}\sum\limits_{t=1}^{T}\left(w\text{log}_{2}\left(1+\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{\text{I}_{t}}\right)-\alpha||{\textbf{D}_{t}}||_{1}\right),$ (17a)
$\text{subject to :}\ d_{mt}=[0,1],\forall m\in\mathcal{M},\ t\in\mathcal{T},$ (17b)
@@
$w\text{log}_{2}\left(1+\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{\text{I}_{t}}\right)\geq\text{SE}^{\text{QoS}}_{t},\forall t\in\mathcal{T},$ (17d)
$\sum\limits_{m=1}^{M}d_{mt}\geq 1,\forall t\in\mathcal{T},$ (17e)
$\text{I}_{t}\geq\sum\limits_{t^{\prime}=1,t^{\prime}\neq t}^{T}\!\!\!\!\!A^{2}p_{u}\eta_{t^{\prime}}^{u}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}(\sum\limits_{m\in\mathcal{M}}\!\!\!d_{mt}\gamma_{mt}\frac{\beta_{mt^{\prime}}}{\beta_{mt}})^{2}+\sum\limits_{t^{\prime}=1}^{T}\!\!\!Ap_{u}\eta_{t^{\prime}}^{u}\sum\limits_{m\in\mathcal{M}}\!\!\!d_{mt}\gamma_{mt}\beta_{mt^{\prime}}+\sum\limits_{m\in\mathcal{M}}\!\!\!Ad_{mt}\gamma_{mt},\forall t\in\mathcal{T},$ (17f)
Considering the non-convexity due to the ratio term inside the log in our objective function (17a), we introduce the lower-bounding auxiliary variable $\Gamma_{t}$. Then, the objective function with an additional constraint as follows :
$\underset{\mathbf{\eta^{u}},\textbf{D},\textbf{I},\mathbf{\Gamma}}{\max~{}}\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right),$ (18a)
$\text{subject to :}\ \Gamma_{t}\leq\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{\text{I}_{t}},\ t\in\mathcal{T},$ (18b)
Taking the partial Lagrangian of the objective function (18a) by considering the constraint (18b) w.r.t. $\mathbf{\Gamma}$. Then :
$L(\mathbf{\eta^{u}},\textbf{D},\textbf{I},\mathbf{\Gamma},\mathbf{\lambda})=\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}-\lambda_{t}\left(\Gamma_{t}-\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{\text{I}_{t}}\right)\right)$ (19)
After computing the stationary point of (19) with respect to $\mathbf{\Gamma}$, we get the optimal value of $\lambda^{*}_{t}=\frac{w^{\prime}}{1+{\Gamma^{*}_{t}}}$. where $w^{\prime}=w/log2$ and from the trivial solution optimal value of $\mathbf{\Gamma}$ is given by :
$\Gamma^{*}_{t}=\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{\text{I}_{t}}$ (20)
Substituting this $\lambda^{*}_{t}$ in (19) and using the Lagrangian dual transformation [16] then new optimization problem can be formulated as :
$\underset{\mathbf{\eta^{u}},\textbf{D},\textbf{I},\mathbf{\Gamma}}{\max~{}}\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}-\left(\Gamma_{t}-\frac{w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+\text{I}_{t}}\right)\right),$ (21a)
@@
It can be observed that after substituting the optimal value of $\mathbf{\Gamma}$, as in (20), in equation (21a), the optimization problem (III) is exactly equal to solving the (III).
If we look at the optimization problem (III), we still have some non-linearities and non-convexities due to product of two variables as well as ratio term in the objective function. Firstly, we tried to handle the ratio term in the objective function using the quadratic transformation approach [16]. The quadratic transformation is performed on the objective function (21a) such that it becomes :
$\underset{\mathbf{\eta^{u}},\textbf{D},\textbf{I},\mathbf{\Gamma},\textbf{u}}{\max~{}}\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right)\\ -\sum\limits_{t=1}^{T}\left(\Gamma_{t}-2u_{t}\sqrt{w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}+u_{t}^{2}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+\text{I}_{t}}\right),$ (22)
The quadratic transformation $2u_{t}\sqrt{w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}-u_{t}^{2}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+\text{I}_{t}}$ is the lower bound and concave approximation of ratio term $\frac{w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+\text{I}_{t}}$ with respect to parameter $u_{t}$.The optimal value of u can be calculated from Equation (23). So the equation (22) is the lower bound to the equation (21a), thus making the approximation feasible.
$u^{*}_{t}\overset{\Delta}{=}\frac{\sqrt{w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+\text{I}_{t}},$ (23)
Our final optimization problem becomes:
@@
After substituting the optimal value of u in the Equation (22), then solving optimization problem (III) is equivalent to the solving of (III).
Notice that our problem (III) is a convex optimization problem in different blocks. When fixing the D, $\mathbf{\Gamma}$ and u, then objective function of problem (24a) becomes concave w.r.t. $\mathbf{\eta^{u}}$ and I. Then, the convex-optimization problem becomes:
@@
Similarly, upon fixing the $\mathbf{\eta^{u}}$, $\mathbf{\Gamma}$ and u, the objective function of (III) becomes concave w.r.t D and I. The convex optimization problem becomes :
@@
So the problem can be solved following the proposed iterative approach. The detailed stepsof the proposed approach are given in Algorithm 1.
1:  Feasible Initialization: $\mathbf{\eta_{k}^{u}}^{(0)}$, $\textbf{D}^{(0)}$ and $i=0$, $\forall k$.
2:  Compute $I_{t}^{(0)}$ from the equation (14) respectively.
3:  repeat
4:     $i=i+1$.
5:     for all t do
6:        Compute $\Gamma_{t}$ using (20) and $u_{t}$ using (23).
7:     end for
8:     Solve optimization problem in (III) for $\mathbf{\eta_{k}^{u}}^{(i)}$ and $\textbf{I}^{(i)}$.
9:     for all t do
10:        Compute again $\Gamma_{t}$ using (20) and $u_{t}$ again using (23).
11:     end for
12:     Solve optimization problem in (III) for $\textbf{D}^{(i)}$ and $\textbf{I}^{(i)}$.
13:     Update ${\mathbf{\eta^{u}}}^{(i+1)}={\mathbf{\eta^{u}}}^{(i)}$, $\textbf{D}^{(i+1)}=\textbf{D}^{(i)}$ and $\textbf{I}^{(i+1)}=\textbf{I}^{(i)}$.
14:  until Convergence
@@
The convergence proof of the proposed Algorithm 1 is given in Appendix V.
@@ IV Performance Evaluation
We performed numerical simulations on a $1000\times 1000$ square meter geographical area with APs and user equipment UEs randomly distributed. To simulate an infinite area network, a wrapped-around approach is used [6]. Large-scale fading coefficients are modeled using a three-slope approach for path loss, incorporating an 8 dB standard deviation for shadow fading. The minimum QoS parameter is taken as 0.1 bits/s/Hz for each UE. Throughout our simulations, we keep $T=30$, $M=50\ \text{or}\ 100$, $A=4$, $\alpha=0.0005\ \text{or}\ \alpha=0.001$, $L_{c}=200$ and $L_{p}=5$, while other parameters stay exactly the same as those outlined in [3] or as explicitly stated. All these numerical results are averaged over $500$ simulations.
@@
Figure 1: The variation of maximum fronthaul load on a AP, objective function and sum SE, when $M=50$.
Figure 2: The variation of maximum front-haul load on a AP, objective function and sum SE, when $M=100$.
@@
Figure 1 shows the variations in maximum front-haul load [^1] The computation of front-haul load in our framework is expressed mathematically as $\sum_{t\in\mathcal{T}}d_{mt}\text{SE}^{u}_{t}\ \forall m$. Additionally, the maximum front-haul is defined as the highest front-haul load among all APs. This metric provides insight into the peak load on the front-haul infrastructure across the entire network., objective function, and sum SE for two different regularization coefficient values, with the number of antennas fixed at $50$. This highlights how the penalty function affects the network’s maximum front-haul load and total sum-throughput. Notably, the objective function value consistently remains lower than the total SE value, exhibiting a drop attributed to the penalty function. When the regularization coefficient is increased by two times, the maximum front-haul load on an AP decreases by $14\%$. This reduction is due to the increased regularization coefficient, which limits AP-UE associations. As a consequence of the lesser number of AP-UE associations, there is a reduction in sum SE by nearly $1\%$. Therefore, the proposed approach effectively decreases the maximum front-haul load while not significantly reducing overall sum throughput, and this favorable outcome corresponds to careful tuning of regularization coefficient.
Similarly, Figure 2 illustrates variability in maximum front-haul load, objective function, and sum SE for two different regularization coefficient values, with $M$ set to 100. For this scenario, when the regularization coefficient is increased two times, the maximum front-haul load and total SE show a similar trend of decrement same as in Fig. 1.
@@
Examining Figures 1 and 2, we observe a noteworthy trend as the number of antennas ($M$) rises from 50 to 100. The maximum front-haul load increases by $14\%$ and $17\%$ with regularization coefficients ($\alpha$) of 0.0005 and 0.001, respectively. This growth can be attributed to the growing number of APs in the fixed area. As the AP count increases, more APs become closer to the each UE, resulting in a higher percentage of APs offering a strong channel or a larger large-scale fading coefficient for each UE. This increasing influence of numerous APs on the large-scale fading coefficient causes a change in the influence of the penalty function, allowing for the association of more APs with each UE. While this increased association raises the maximum front-haul load, it is important to note that increasing the number of APs results in a significant $30\%$ boost in overall sum-throughput. In essence, the system responds to the increased AP count by increasing the strength of channels across numerous APs, hence increasing the network’s overall throughput. The penalty function is a dynamic component of this process, optimizing AP-UE associations to optimize the objective function while reducing negative affects via penalties.
Figure 3: The variation of sum SE for different scenarios, when $\alpha=0.001$.
@@
Indeed, our proposed technique aims to enhance the sum SE. Figure 4 shows how our strategy has a significant influence on the $90\%-likely$ per-user SE. Our proposed scheme outperforms scenarios ’a’, ’b’, ’c’, and ’d’ in terms of $90\%-likely$ per-user SE by $175\%$, $22\%$, $4.7\%$, and $11.7\%$, respectively, at $M=50$. Similarly, for $M=100$, our approach outperforms the same cases by $52.3\%$, $6.7\%$, $2.1\%$, and $3.2\%$ in terms of $90\%-likely$ per-user SE. The findings illustrate an additional benefit of the method we propose, demonstrating its massive contribution to improving per-user throughput while also increasing total throughput.
@@
Figure 5, shows the variation of SE with and without the discretization of matrix D. The figure shows that for $m=50$, the variation of SE is negligible with and without the discretization of elements of matrix D. For the case when $m=100$, the difference in SE is mere $0.2\%$. The negligible SE variation can be attributed the fact that penalty parameter does not force elements of matrix D towards exact zero or one. This highlights the importance of penalty function in successfully forcing the value of elements of matrix D almost towards zero or one, thus our proposed solution does not require additional algorithm for AP-UE association.
@@
The convergence trajectory of Algorithm 1 is portrayed in Figure 6. The figure shows that as the number of iterations increases, so does the value of the objective function, leading up in a non-decreasing function. However, the rate of rise gradually decreases, ensuring that the rate of increase of objective function finally becomes lower than the defined tolerance of $5\times 10^{-3}$. This observation guarantees the convergence of Algorithm 1.

### 26 Feb 2024 v2

@@ III Proposed Joint Solution using Relaxation Technique
Problem (II-C) is a mixed-integer type problem which has a very high complexity, so we first convert the discrete binary constraint (4b) into a continuous one, resulting in:
$d_{mt}\in[0,1],\forall m\in\mathcal{M},\ t\in\mathcal{T}.$ (5)
Remark 1 : The introduction of l1-penalty in the objective function serves to induce the sparsity in the AP-UE association matrix D. By careful selection of the trade-off factor $\alpha$, controlled-sparsity is achieved by limiting the number of APs that serve the UEs. Therefore, the matrix D has non-zero values only for the chosen AP-UE pairs. After completion of AP-UE association, all entries in D undergo rounding to the nearest integer values i.e., zero and one, thus maintaining the feasibility of the optimization problem (4a). In the forthcoming results section, the SE with and without rounding-off of values in D are compared to numerically show the effectiveness of our proposed approach.
Then, the optimization problem (4a) can be formulated as:
$\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}$ $\ \ \sum\limits_{t=1}^{T}\left(w\text{log}_{2}\left(\!1+\Gamma^{*}_{t}\right)-\alpha||{\textbf{D}_{t}}||_{1}\right),$ (6a)
subject to : $\ \ d_{mt}\in[0,1],\forall m\in\mathcal{M},\ t\in\mathcal{T},$ (6b)
@@
$\ \ w\text{log}_{2}\left(1+\Gamma^{*}_{t}\right)\geq\text{SE}^{\text{QoS}}_{t},\forall t\in\mathcal{T},$ (6d)
$\ \ \sum\limits_{m=1}^{M}d_{mt}\geq 1,\forall t\in\mathcal{T}.$ (6e)
Considering the non-convexity due to the ratio term inside the log in our objective function (6a), we introduce the lower-bounding auxiliary variable $\Gamma_{t}$, resulting in the objective function with an additional constraint as follows:
$\underset{\bm{\eta^{u}},\textbf{D},\mathbf{\Gamma}}{\max~{}}$ $\ \ \sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right),$ (7a)
subject to : $\ \ \Gamma_{t}\leq\Gamma^{*}_{t},\ t\in\mathcal{T},$ (7b)
where $\mathbf{\Gamma}=[\Gamma_{1},\Gamma_{2},...\Gamma_{T}]^{\intercal}$. Taking the partial Lagrangian of objective function (7a) by considering constraint (7b) with respect to $\mathbf{\Gamma}$, we have:
$\begin{split}L(\bm{\eta^{u}},\textbf{D},\mathbf{\Gamma},\bm{\lambda})=&\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right)\\ &-\sum\limits_{t=1}^{T}\lambda_{t}\left(\Gamma_{t}-\Gamma^{*}_{t}\right),\end{split}$ (8)
where $\bm{\lambda}=[\lambda_{1},\lambda_{2},...\lambda_{T}]^{\intercal}$.
After computing the stationary point of (8) with respect to $\mathbf{\Gamma}$, we get the optimal value of $\lambda^{*}_{t}=\frac{w^{\prime}}{1+{\Gamma^{*}_{t}}}$. where $w^{\prime}=w/{\ln 2}$ with the optimal value of $\mathbf{\Gamma}=\mathbf{\Gamma^{*}}$, where $\mathbf{\Gamma^{*}}=[\Gamma^{*}_{1},\Gamma^{*}_{2},...\Gamma^{*}_{T}]^{\intercal}$. Substituting this $\lambda^{*}_{t}$ in (8) and using the Lagrangian dual transformation [17], the problem in (III) can be reformulated as :
$\begin{split}&\underset{\bm{\eta^{u}},\textbf{D},\mathbf{\Gamma}}{\max~{}}\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right)-\sum\limits_{t=1}^{T}\left(\Gamma_{t}-\frac{w^{\prime}(1+{\Gamma}_{t})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+{I}_{t}}\right),\\ \end{split}$ (9a)
@@
It can be observed that after substituting the optimal value of $\mathbf{\Gamma}$, as in (2), in equation (9a), the optimization problem in (III) reduces exactly to the one in (III). If we look at optimization problem (III), we still have some non-linearities and non-convexities due to a ratio term and a product of two variables in the objective function. First, we attempt to handle the ratio term in the objective function using the quadratic transformation approach [17]. The quadratic transformation is performed on the objective function (9a) such that it becomes:
$\begin{split}\underset{\bm{\eta^{u}},\textbf{D},\mathbf{\Gamma},\textbf{u}}{\max~{}}&\ \ \sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right)-\sum\limits_{t=1}^{T}\left(\Gamma_{t}-2u_{t}\sqrt{w^{\prime}(1+{\Gamma}_{t})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}\right)\\ &+\sum\limits_{t=1}^{T}u_{t}^{2}\left({A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+{I}_{t}}\right),\end{split}$ (10)
where $\textbf{u}=[u_{1},u_{2},...u_{T}]^{\intercal}$.
The quadratic transformation in equation (10) is the lower bound and concave approximation of ratio term $\frac{w^{\prime}(1+{\Gamma}_{t})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+{I}_{t}}$ with respect to the parameter $u_{t}$. The optimal value of u can be calculated from equation (11).
$u^{*}_{t}\overset{\Delta}{=}\frac{\sqrt{w^{\prime}(1+{\Gamma}_{t})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+{I}_{t}}.$ (11)
Therefore, equation (10) is the lower bound to equation (9a), thus making the approximation feasible.
The optimization problem becomes:
@@
After substituting the optimal value of u in equation (10) and then solving optimization problem (III) is equivalent to solving Problem (III). Notice that Problem (III) is a convex optimization problem in different blocks. If we fix D, $\mathbf{\Gamma}$ and u, then the objective function of problem (12a) becomes concave with respect to $\bm{\eta^{u}}$. Then, the problem reduces to:
@@
Similarly, upon fixing $\bm{\eta^{u}}$, $\mathbf{\Gamma}$ and u, the objective function of (III) becomes concave with respect to D. Then, the optimization problem becomes:
@@
Problem III can be solved following the proposed iterative approach as follows. This problem is decomposed into two distinct blocks, denoted as Problem III and Problem III. Initially, the solution for Problem III is obtained while keeping D, $\mathbf{\Gamma}$, and u fixed. Subsequently, based on the solution derived for Problem III, the solution for Problem III is determined. This is repeated until the convergence criterion is met. Algorithm 1 provides the details of this approach.
1:  Feasible Initialization: ${\bm{\eta_{k}^{u}}}^{(0)}$, $\textbf{D}^{(0)}$ and $i=0$, $\forall k$.
2:  repeat
3:     $i=i+1$.
4:     for all t do
5:        Compute $\Gamma_{t}^{(i)}$ using (2) and $u_{t}^{(i)}$ using (11).
6:     end for
7:     Solve optimization problem in (III) for ${\bm{\eta_{k}^{u}}}^{(i)}$.
8:     for all t do
9:        Again compute $\Gamma_{t}^{(i)}$ using (2) and $u_{t}^{(i)}$ using (11).
10:     end for
11:     Solve optimization problem in (III) for $\textbf{D}^{(i)}$.
12:     Update ${\bm{\eta^{u}}}^{(i+1)}={\bm{\eta^{u}}}^{(i)}$ and $\textbf{D}^{(i+1)}=\textbf{D}^{(i)}$.
13:  until $\left|\frac{f_{2}^{(i+1)}-f_{2}^{(i)}}{f_{2}^{(i)}}\right|\leq\epsilon$
@@
For the convergence proof of the proposed Algorithm 1, please refer to the appendixV.
@@ IV Performance Evaluation
We performed numerical simulations on a $1000\times 1000$ square meters geographical area with APs and user equipment UEs randomly distributed. To simulate an infinite area network, a wrapped-around approach is used [6]. Large-scale fading coefficients are modeled using a three-slope approach for path loss, incorporating an 8 dB standard deviation for shadow fading. The minimum QoS parameter is taken as 0.1 bits/s/Hz for each UE. Throughout our simulations, we keep $T=30$, $M=50\ \text{or}\ 100$, $A=4$, $\alpha=0.0005\ \text{or}\ \alpha=0.001$, $L_{c}=200$ and $L_{p}=5$, while other parameters stay exactly the same as those outlined in [3] or as explicitly stated. All these numerical results are averaged over $500$ simulations.
@@
Figure 1: The variation of maximum fronthaul load on an AP, objective function, and the sum SE, when $M=50$.
@@
Figure 1 shows the variations in maximum front-haul load, objective function, and the sum SE for two different regularization coefficient values, with the number of antennas fixed at $50$. This highlights how the penalty function affects the network’s maximum front-haul load and total sum-throughput. Notably, the objective function value consistently remains lower than the total SE value, exhibiting a drop attributed to the penalty function. When the regularization coefficient is increased by two times, the maximum front-haul load on an AP decreases by $14\%$. This reduction is due to the increased regularization coefficient, which limits AP-UE associations. As a consequence of the lesser number of AP-UE associations, there is a reduction in the sum SE by nearly $1\%$. Similarly, Figure 2 illustrates variability in maximum front-haul load, objective function, and sum SE for two different regularization coefficient values, however with $M$ set to 100. For this scenario too, when the regularization coefficient is increased two times, the maximum front-haul load and total SE show a similar decreasing trend as in Fig. 1. Therefore, the proposed approach effectively decreases the maximum front-haul load while not significantly reducing overall sum throughput, and this is achieved by a careful tuning of regularization coefficient.
Figure 2: The variation of maximum front-haul load on an AP, objective function, and the sum SE, when $M=100$.
@@
Examining Figures 1 and 2, we observe a noteworthy trend as the number of antennas ($M$) increases from 50 to 100. The maximum front-haul load increases by $14\%$ and $17\%$ with regularization coefficients ($\alpha$) of 0.0005 and 0.001, respectively. This growth can be attributed to the growing number of APs in the fixed area. As the AP count increases, there are more APs in the vicinity of each UE, resulting in a higher percentage of APs offering a strong channel or a larger large-scale fading coefficient for each UE. This gainful influence of numerous APs on the large-scale fading coefficient causes a change in the influence of the penalty function, allowing more APs to associate with each UE. While this increased association raises the maximum front-haul load, it also results in a significant ($30\%$) boost in the overall sum-throughput. In essence, the system responds to the increased AP count by increasing the strength of channels across numerous APs, hence increasing the network’s overall throughput. The penalty function is a dynamic component of this process, optimizing AP-UE associations to optimize the objective function while reducing negative affects via penalties.
Figure 3: The variation of the sum SE for different scenarios, when $\alpha=0.001$.
@@
Indeed, our proposed technique aims to enhance the sum SE. Figure 4 shows how our strategy has a significant influence on the $90\%-likely$ per-user SE. Our proposed scheme outperforms scenarios ’a’, ’b’, ’c’, and ’d’ in terms of $90\%-likely$ per-user SE by $175\%$, $22\%$, $4.7\%$, and $11.7\%$, respectively, at $M=50$. Similarly, for $M=100$, our approach outperforms the same cases by $52.3\%$, $6.7\%$, $2.1\%$, and $3.2\%$ in terms of $90\%-likely$ per-user SE. The findings illustrate an additional benefit of the method we propose, demonstrating its massive contribution to improve per-user throughput while also increasing the total throughput.
@@
In Figure 5, we provide a numerical justification for relaxation proposed in (5). It shows the variation of the SE with and without the discretization of matrix D. The figure shows that for $m=50$, the variation in the SE is negligible with and without the discretization of elements of matrix D. For the case when $m=100$, the difference in SE is mere $0.2\%$. The negligible SE variation can be attributed to the fact that penalty parameter does not force elements of matrix D too much away from zero or one, thus our proposed solution does not require additional algorithm for AP-UE association.
@@
In Figure 6, we provide a numerical evidence for the convergence Algorithm 1. The figure shows that as the number of iterations increases, so does the value of the objective function, leading up in a non-decreasing function. However, the rate of rise gradually decreases, ensuring that the rate of increase of objective function finally becomes lower than the defined tolerance of $\epsilon=5\times 10^{-3}$. This observation guarantees the convergence of Algorithm 1.